# The calculus of Duistermaat’s triple index

### v2

@@ 2. Existence of the Duistermaat index
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})=n_{-}\Big(R_{2}^{\epsilon}-R_{1}^{\epsilon}\Big),\qquad\text{for all }0<\epsilon\ll 1.$ (2.21)
Equation (2.21) was already established in [5, Eq. (3.32)], but we repeat the proof here for completeness. We use equation (2.7) with $R^{\epsilon}_{3}=\epsilon^{-1}I$, and the desired result follows once we establish that $n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)=0=n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)$. In other words, we claim that for any Lagrangian frame $(X,Y)^{T}$, the Hermitian operator $\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}$ is non-negative definite for small enough $\epsilon>0$. It is equivalent to consider
@@ 3.1. General identities
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})$ $=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}$ (3.2)
@@
    $=\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{1})-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}$ (3.6)
@@
    (Anti-symplectic transformation) If $\mathsf{A}$ is an anti-symplectic linear transformation on $(\mathbb{C}^{n}\oplus\mathbb{C}^{n},\omega)$, i.e., a bijection that satisfies $\omega(\mathsf{A}x,\mathsf{A}y)=-\omega(x,y)$, then
@@
The first two equalities in (3.1) follow directly from (2.7); for the third we also use (2.17) and (2).
@@
Symplectic additivity (3.7) follows by noting that a possible choice of frame for $\mathcal{L}\oplus\mathcal{M}$ is the direct sum of frames. Therefore, for all but finitely many $\epsilon$, one has $R^{\epsilon}_{\mathcal{L}\oplus\mathcal{M}}=R^{\epsilon}_{\mathcal{L}}\oplus R^{\epsilon}_{\mathcal{M}}$. Equation (3.7) then follows from (2.7) and additivity of $n_{-}$ under direct sums.
@@ 3.2. Self-adjoint linear relations
For a given splitting of the symplectic space (in our case, $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$), every Lagrangian plane can be viewed as a self-adjoint linear relation; see [22, Sec. 14.1] or [8, Sec. 4.2]. In this context, the difference of two Lagrangian planes, $\mathcal{L}$ and $\mathcal{M}$, is the Lagrangian plane
$\mathcal{L}-\mathcal{M}:=\left\{\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}-v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}\end{pmatrix}\in\mathcal{L},\ \begin{pmatrix}u\\ v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathcal{M}\right\}$, (3.11)
and the inverse of a Lagrangian plane $\mathcal{L}$ is
$\mathcal{L}^{-1}:=\left\{\begin{pmatrix}v\\ u\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\right\}$. (3.12)
If we define
$\operatorname{dom}\mathcal{L}$ $:=\left\{u\in\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\text{ for some }v\in\mathbb{C}^{n}\right\},$ (3.13)
$\operatorname{mul}\mathcal{L}$ $:=\left\{v\in\mathbb{C}^{n}:\begin{pmatrix}0\\ v\end{pmatrix}\in\mathcal{L}\right\}$, (3.14)
we have $\operatorname{mul}\mathcal{L}=\left(\operatorname{dom}\mathcal{L}\right)^{\perp}$ and $\mathcal{L}$ can be represented as
@@
for some Hermitian matrix $L\colon\operatorname{dom}\mathcal{L}\to\operatorname{dom}\mathcal{L}$, see, e.g. [3, Thm. 1.5.1] or [22, Prop. 14.2].
@@
Remark 3.3.
It is tempting to drop the condition that the first argument in (3.18) is a graph and conjecture that $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},{\mathcal{G}_{\infty}})=\mathfrak{n}_{-}(\mathcal{L}_{2}-\mathcal{L}_{1})$ for any Lagrangian planes $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, where the difference $\mathcal{L}_{2}-\mathcal{L}_{1}$ is computed using (3.11) and the negative index $\mathfrak{n}_{-}$ of a Lagrangian plane $\mathcal{L}$ could be defined as the Morse index of its operator part $L$, see (3.15). This enticing conjecture is unfortunately false, as evidenced by the extra term in (3.19).
@@
For Part (3) we observe that
@@
$=\big\{(x,(L-PA)x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big\}.$
Using (3.16), we reduce equations (3.18) and (3.19) to their $A=0$ versions,
$\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})=n_{-}(L),\qquad\operatorname{\iota}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})=n_{-}(-L)+\dim\mathcal{L}\cap\mathcal{V}.$ (3.23)
@@
$\mathbb{C}^{n}\oplus\mathbb{C}^{n}=\big(\operatorname{dom}\mathcal{L}\oplus\operatorname{dom}\mathcal{L}\big)\oplus\big(\operatorname{mul}\mathcal{L}\oplus\operatorname{mul}\mathcal{L}\big).$ (3.24)
It is straightforward to check that $\omega$ restricted to each space is symplectic and the sum of the restrictions is the original $\omega$. With respect to this partition (and with some abuse of notation), we have ${\mathcal{G}_{0}}={\mathcal{G}_{0}}\oplus{\mathcal{G}_{0}}$, $\mathcal{L}=\mathcal{G}_{L}\oplus{\mathcal{G}_{\infty}}$ and ${\mathcal{G}_{\infty}}={\mathcal{G}_{\infty}}\oplus{\mathcal{G}_{\infty}}$. We then use symplectic additivity (3.7), as well as the special values (3.1), to compute
$\operatorname{\iota}_{\mathbb{C}^{n}}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}({\mathcal{G}_{0}},\mathcal{G}_{L},{\mathcal{G}_{\infty}})+\operatorname{\iota}_{\operatorname{mul}\mathcal{L}}({\mathcal{G}_{0}},{\mathcal{G}_{\infty}},{\mathcal{G}_{\infty}})$ (3.25)
$=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}({\mathcal{G}_{0}},\mathcal{G}_{L},{\mathcal{G}_{\infty}})=n_{-}(L),$
$\operatorname{\iota}_{\mathbb{C}^{n}}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}(\mathcal{G}_{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})+\operatorname{\iota}_{\operatorname{mul}\mathcal{L}}({\mathcal{G}_{\infty}},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})$ (3.26)
$=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}(\mathcal{G}_{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})+\dim\operatorname{mul}\mathcal{L}=n_{-}(-L)+\dim\mathcal{L}\cap{\mathcal{G}_{\infty}},$
@@ 4.1. The Kashiwara index
Another commonly used index of a triple of Lagrangian planes arose independently in the work of Hörmander, Kashiwara and Wall; see, for instance, [23], [15, Eq. (3.3.15)], [17, Def. 1.5.1], [9, Sec. 8], [12, Sec 4.2]. The Hörmander–Kashiwara–Wall index $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is defined [^4] Only the latter reference considered complex symplectic space; the signs in our definition of the form $\Omega$ are different. as the signature of the form $\Omega$ from Theorem 2.1,
@@
The advantage in dealing with $\operatorname{Sig}(A)$ is its symmetry property $\operatorname{Sig}(-A)=-\operatorname{Sig}(A)$, in contrast to (2.17). However, our motivation comes, ultimately, from operators that are bounded only from below; in this case $n_{-}$ may be well-defined, while $n_{+}$ will not be.
@@
$\dim\mathcal{L}_{j}\cap\mathcal{L}_{k}=\dim\mathcal{M}_{j}\cap\mathcal{M}_{k},\quad\text{for all }j,k\in\{1,2,3\},$ (4.5)
@@
Proposition 4.2 is well-known when (4.7) is replaced by the condition $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$, see, e.g. [20, Thm. 8.4(d)] or [1, Prop. 4.4]. But the indices $s$ and $\operatorname{\iota}$ are related by an invertible linear transformation involving the invariants appearing in (4.5). ∎
Corollary 4.3.
For any two pairs $(\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{2})$ of Lagrangian planes, there is a linear symplectic transformation $\mathsf{S}$ satisfying $\mathsf{S}\mathcal{L}_{j}=\mathcal{M}_{j}$, $j=1,2$, if and only if
$\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}=\dim\mathcal{M}_{1}\cap\mathcal{M}_{2}.$ (4.8)
Proof.
While usually established before the more complicated case of a triple, see [20, Exercise 8.1], this fact also follows from Proposition 4.2 by applying it to the triples $(\mathcal{L}_{1},\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{1},\mathcal{M}_{2})$ and using (3.1). ∎

### v3

@@ 2.1. Existence of the index
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})=n_{-}\Big(R_{2}^{\epsilon}-R_{1}^{\epsilon}\Big)\qquad\text{for all }0<\epsilon\ll 1.$ (2.22)
Equation (2.22) was already established in [15, Eq. (3.32)], but we repeat the proof here for completeness. We use equation (2.7) with $R^{\epsilon}_{3}=\epsilon^{-1}I$, and the desired result follows once we establish that $n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)=0=n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)$. In other words, we claim that for any Lagrangian frame $(X,Y)^{T}$, the Hermitian operator $\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}$ is non-negative definite for small enough $\epsilon>0$. It is equivalent to consider
@@
2.2. Proofs of Theorems 1.1 and 1.4
Proof of Theorem 1.1.
Theorem 2.1 proves the existence of a function $\operatorname{\iota}$ satisfying the axioms in Theorem 1.1. Uniqueness follows from the observation that these axioms provide a prescription for the computation of $\operatorname{\iota}$. First we reduce the computation to the case when $\mathcal{L}_{3}$ is transversal to both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ by using the cocycle identity with some $\mathcal{L}_{4}$ that is transversal to the other three. We can then find a symplectic transformation that maps $\mathcal{L}_{3}$ to $\mathcal{G}_{\infty}$} and $\mathcal{L}_{1}$ to $\mathcal{G}_{0}$} (which is well-known to be possible, see for example [46, Exercise 8.1]). Since this transformation preserves transversality, the transformed $\mathcal{L}_{2}$ can be represented as the graph of a Hermitian matrix, and (1.6) can be used. ∎
Proof of Theorem 1.4.
Choosing the frame $(I,A)$ for the Lagrangian plane $\mathcal{L}_{1}=\mathcal{G}_{A}$ and setting $\epsilon=0$, we get $R_{1}^{0}=A$. The same process works for $\mathcal{G}_{B}$ and $\mathcal{G}_{C}$, and (1.10) follows from (2.7). ∎
@@ 3.1. General identities
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})$ $=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2},$ (3.2)
@@
    $=\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{1})-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}.$ (3.6)
@@
    (Anti-symplectic transformation) If $\mathsf{A}$ is an anti-symplectic linear transformation on $(\mathbb{C}^{n}\oplus\mathbb{C}^{n},\omega)$, i.e., a bijection that satisfies $\omega(\mathsf{A}u,\mathsf{A}v)=-\omega(u,v)$, then
@@
The first two equalities in (3.1) follow directly from (2.7); for the third we also use (2.18) and (2.1).
@@
Symplectic additivity (3.7) follows by noting that a possible choice of frame for $\mathcal{L}\oplus\mathcal{M}$ is the direct sum of frames. Therefore, for all but finitely many $\epsilon$, one has $R^{\epsilon}_{\mathcal{L}\oplus\mathcal{M}}=R^{\epsilon}_{\mathcal{L}}\oplus R^{\epsilon}_{\mathcal{M}}$. Equation (3.7) then follows from (2.7) and the additivity of $n_{-}$ under direct sums.
@@ 3.2. Self-adjoint linear relations
For a given splitting of the symplectic space (in our case $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$), every Lagrangian plane can be viewed as a self-adjoint linear relation; see [49, Sec. 14.1] or [17, Sec. 4.2]. In this context, the difference of two Lagrangian planes, $\mathcal{L}$ and $\mathcal{M}$, is the Lagrangian plane
$\mathcal{L}-\mathcal{M}:=\left\{(x,y_{{}_{\mathcal{L}}}-y_{{}_{\mathcal{M}}})\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:(x,y_{{}_{\mathcal{L}}})\in\mathcal{L},\ (x,y_{{}_{\mathcal{M}}})\in\mathcal{M}\right\}.$ (3.11)
This definition is chosen so that
$\mathcal{G}_{A}-\mathcal{G}_{B}=\mathcal{G}_{A-B}$ (3.12)
for any $n\times n$ Hermitian matrices $A$ and $B$. Another important special case is
$\mathcal{L}-{\mathcal{G}_{\infty}}={\mathcal{G}_{\infty}}={\mathcal{G}_{\infty}}$-$\mathcal{L}$ (3.13)
for any Lagrangian plane $\mathcal{L}$.
The inverse of a Lagrangian plane $\mathcal{L}$ is
$\mathcal{L}^{-1}:=\left\{(y,x)\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:(x,y)\in\mathcal{L}\right\}.$ (3.14)
We define by
$\operatorname{dom}\mathcal{L}$ $:=\left\{x\in\mathbb{C}^{n}:(x,y)\in\mathcal{L}\text{ for some }y\in\mathbb{C}^{n}\right\}$ (3.15)
$\operatorname{mul}\mathcal{L}$ $:=\left\{y\in\mathbb{C}^{n}:(0,y)\in\mathcal{L}\right\}$ (3.16)
the domain and the multivalued part of the linear relation $\mathcal{L}$. For instance, if $M$ is an $(n\times n)$ matrix then $\operatorname{dom}\mathcal{G}_{M}=\mathbb{C}^{n}$, $\operatorname{mul}\mathcal{G}_{M}=\{0\}$ and $\operatorname{mul}\mathcal{G}_{M}^{-1}=\ker M$; on the other hand $\operatorname{dom}\mathcal{G}_{\infty}=\{0\}$ and $\operatorname{mul}\mathcal{G}_{\infty}=\mathbb{C}^{n}$. In general, we have $\operatorname{mul}\mathcal{L}=\left(\operatorname{dom}\mathcal{L}\right)^{\perp}$ and $\mathcal{L}$ can be represented as
@@
for some Hermitian matrix $L\colon\operatorname{dom}\mathcal{L}\to\operatorname{dom}\mathcal{L}$, as in [12, Thm. 1.5.1] or [49, Prop. 14.2].
@@
For Part (3) we use (3.18) to get
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{L},{\mathcal{G}_{\infty}})=\operatorname{\iota}(\mathcal{G}_{0},\mathcal{L}-\mathcal{G}_{A},{\mathcal{G}_{\infty}}),\qquad\operatorname{\iota}(\mathcal{L},\mathcal{G}_{A},{\mathcal{G}_{\infty}})=\operatorname{\iota}(\mathcal{L}-\mathcal{G}_{A},\mathcal{G}_{0},{\mathcal{G}_{\infty}}).$ (3.24)
We observe that
@@
$=\big\{(x,(L-PA)x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big\}$
$=\big\{(x,(L-A_{\operatorname{dom}\mathcal{L}})x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big\}.$
This shows that it is enough to establish the $A=0$ versions of (3.20) and (3.21), namely
$\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})=n_{-}(L),\qquad\operatorname{\iota}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})=n_{-}(-L)+\dim\mathcal{L}\cap\mathcal{G}_{\infty}.$ (3.26)
@@
$\mathbb{C}^{n}\oplus\mathbb{C}^{n}=\big(\operatorname{dom}\mathcal{L}\oplus\operatorname{dom}\mathcal{L}\big)\oplus\big(\operatorname{mul}\mathcal{L}\oplus\operatorname{mul}\mathcal{L}\big)=:V^{\prime}\oplus V^{\prime\prime}.$ (3.27)
It is straightforward to check that the spaces $V^{\prime}$ and $V^{\prime\prime}$ are symplectic [^2] The restriction of $\omega$ to $V^{\prime}$ is the canonical symplectic form induced by the inner product on $\operatorname{dom}\mathcal{L}$ (cf. (1.1)), and likewise for the restriction of $\omega$ to $V^{\prime\prime}$. with the forms $\omega^{\prime}:=\omega|_{V^{\prime}}$ and $\omega^{\prime\prime}:=\omega|_{V^{\prime\prime}}$, and that $\omega=\omega^{\prime}\oplus\omega^{\prime\prime}$. With respect to this partition, we have ${\mathcal{G}_{0}}={\mathcal{G}_{0}^{\prime}}\oplus{\mathcal{G}_{0}^{\prime\prime}}$, $\mathcal{L}={\mathcal{G}_{L}^{\prime}}\oplus{\mathcal{G}_{\infty}^{\prime\prime}}$ and ${\mathcal{G}_{\infty}}={\mathcal{G}_{\infty}^{\prime}}\oplus{\mathcal{G}_{\infty}^{\prime\prime}}$. We then use symplectic additivity (3.7), as well as the special values (3.1) and normalization (1.6), to compute
$\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{V^{\prime}}({\mathcal{G}_{0}^{\prime}},{\mathcal{G}_{L}^{\prime}},{\mathcal{G}_{\infty}^{\prime}})+\operatorname{\iota}_{V^{\prime\prime}}({\mathcal{G}_{0}^{\prime\prime}},{\mathcal{G}_{\infty}^{\prime\prime}},{\mathcal{G}_{\infty}^{\prime\prime}})$ (3.28)
$=\operatorname{\iota}_{V^{\prime}}({\mathcal{G}_{0}^{\prime}},{\mathcal{G}_{L}^{\prime}},{\mathcal{G}_{\infty}^{\prime}})=n_{-}(L).$
Similarly, using the modified normalization $\operatorname{\iota}(\mathcal{G}_{A},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})=\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{-A},{\mathcal{G}_{\infty}})=n_{-}(-A)$, obtained from (1.6) and (3.18), we obtain
$\operatorname{\iota}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{V^{\prime}}({\mathcal{G}_{L}^{\prime}},{\mathcal{G}_{0}^{\prime}},{\mathcal{G}_{\infty}^{\prime}})+\operatorname{\iota}_{V^{\prime\prime}}({\mathcal{G}_{\infty}^{\prime\prime}},{\mathcal{G}_{0}^{\prime\prime}},{\mathcal{G}_{\infty}^{\prime\prime}})$ (3.29)
$=\operatorname{\iota}_{V^{\prime}}({\mathcal{G}_{L}^{\prime}},{\mathcal{G}_{0}^{\prime}},{\mathcal{G}_{\infty}^{\prime}})+\frac{1}{2}\dim V^{\prime\prime}=n_{-}(-L)+\dim\mathcal{L}\cap{\mathcal{G}_{\infty}},$
@@ 4.1. The Kashiwara index
Another commonly used index of a triple of Lagrangian planes arose independently in the work of Hörmander, Kashiwara and Wall; see, for instance, [50], [31, Eq. (3.3.15)], [43, Def. 1.5.1], [19, Sec. 8], [26, Sec 4.2]. The Hörmander–Kashiwara–Wall index $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is defined [^3] Only the reference [26] considered complex spaces; the signs in our definition of the form $\Omega$ are different. as the signature of the form $\Omega$ from Theorem 2.1,
@@
The advantage of dealing with $\operatorname{Sig}(A)$ is its symmetry property $\operatorname{Sig}(-A)=-\operatorname{Sig}(A)$, in contrast to (2.18). However, our motivation comes, ultimately, from operators that are bounded only from below; in this case $n_{-}$ may be well-defined while $n_{+}$ will not be.
@@
$\dim\mathcal{L}_{j}\cap\mathcal{L}_{k}=\dim\mathcal{M}_{j}\cap\mathcal{M}_{k}\quad\text{for all }j,k\in\{1,2,3\},$ (4.5)
@@
Proposition 4.2 is well known when (4.7) is replaced by the condition $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$, see [46, Thm. 8.4(d)] or [2, Prop. 4.4]. By Lemma 4.1 the indices $s$ and $\operatorname{\iota}$ are related by an invertible transformation involving the invariants in (4.5) and so, assuming (4.5) holds, $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$ if and only if $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$. ∎